A monopoly firm is faced with the following demand function

P = 26 – 0.5Q. The Marginal Cost function for the firm is given

by 6 + 6Q and the total fixed cost is 4.

Determine;

a) The profit maximizing output.

b) The level of supernormal profit if any.

c) The output level at the break-even point


1
Expert's answer
2021-06-03T13:45:10-0400

"Soln,"

"p=26-0.5Q"

"MC=6+6Q"

"TFC=4"

a. "MC=MR"

P=AR

"P=26-0.5Q"

"TR=P.Q(26-0.5Q)Q"

"TR=26Q-0.5Q^2"

"MR=\frac {\delta TR}{\delta Q}=26-Q"

"MR=26-Q"

Max at MC=MR

"=6+6Q=26-Q= \frac{ 7Q}{7}=\frac {20}{7}"

"Q=2.86 units."


b."TR" greater than TC

"TR=26-0.5Q^2" BUT "Q=2.86"

"26(2.86)-0.5(2.86)^2"

74.36−4.0898=78.4498


"TR=78.4498"

"TC=\int(MC) dQ" BUT "MC=6+6Q"

"TC=4+6Q+6Q^2=4+17.16+49.07=70.23"

"TC=70.23"


"TR-TC"

"=78.4498-70.23\n=8.2122"


c. At break even point TC=TR

"TC=4+6Q+6Q^2"

"TR=26Q-0.5Q^2"

THUS, "26Q-0.5Q^2=4+6Q+6Q^2"

"=6.5Q^2-20Q+4=0"

=42.25-40Q+8=0

"40Q=50.25"

"=1.26"